A culture of E. coli is diluted 1:1000. 1 ml of the 1:1000 dilution is spread onto an agar plate. After incubation, 50 colonies are observed. What is the concentration of the original culture? Show your work and include proper units for full credit.
Answer #1

total dilution used 1:1000 or 1 *10^-3 ; volume plated =1ml , no. of colonies obtained =50

no. of CFU/volume plated (mL) * total dilution used = 50/1 mL* 1*10^-3= 5*10^4 bacteria per mL.